e ^ (- x) * d * y - (e ^ (- x) * y + y ^ 2) * d * x = 0 .
1
Expert's answer
2021-09-23T17:17:09-0400
"e^{-x}dy-(e^{-x}y+y^2)dx=0"

"y'-y=e^xy^2"

"u=y^{1-n}=y^{1-2}=y^{-1}"

"u'=-y^{-2}y'"

"-u'-u=e^x"

"\mu(x)=e^x"

"e^xu'+e^xu=-e^{2x}"

"(e^xu)'=-e^{2x}"

Integrate


"e^xu=\int(-e^{2x})dx"

"e^xu=-\dfrac{1}{2}e^{2x}+C_1"

"u=-\dfrac{1}{2}e^{x}+C_1e^{-x}"

"y(x)=\dfrac{1}{-\dfrac{1}{2}e^{x}+C_1e^{-x}}"

"y(x)=\dfrac{2e^x}{C-e^{2x}}"